If Sally adds four times her age four years from now to five times her age five years from now, she gets ten times her current age. how old will sally be five years from now?
Answer by ankor@dixie-net.com(22740)

If Sally adds four times her age four years from now to five times her age five years from now, she gets ten times her current age.
4(s+4) + 5(s+5) = 10s
4s + 16 + 5s + 25 = 10s
16 + 25 = 10s - 4s - 5s
41 = s
:
Sally is 41 now
:
how old will sally be five years from now?
Sally will be 46